The observations from a random sample of n = 6 from a normal population are:

13.15, 13.72, 12.58, 13.77, 13.01, 13.06.



Test the null hypothesis of H0:μ=13 against the alternative hypothesis of H1:μ<13. Use a 5% level of significance. Answer the following, rounding off your answer to three decimal places.

(a) What is the sample mean?

(b) What is the sample standard deviation?

(c) What is the test statistic used in the decision rule?

(d) Can the null hypothesis be rejected at a 5% significance level?

A. Do not reject the null hypothesis.
B. Reject the null hypothesis.
Answer #1

a)

Sample mean = = X / n = 13.215

b)

Sample standard deviation = sqrt [ X2 - n 2 / ( n - 1) ] = 0.455

c)

Test statistics

t = ( - ) / ( S / sqrt(n) )

= ( 13.215 - 13) / ( 0.455 / sqrt(6) )

= -1.157

d)

df = n - 1 = 6 - 1 = 5

From T table,

t critical value at 0.05 significance level with 5 df = -2.015

Since test statistics > -2.015, Do not reject the null hypothesis